Vapor Liquid Equilibria

Compute the vapor-liquid equilibria for a system of methanol and ethanol at 1 atm and plot the x − y diagram for this system. The following Antoine’s coefficents may be used:

Compound B.Pt.  A  B  C
Methanol 337.8 K 5.20409 1581.341 -33.50
Ethanol  351.5 K 5.24677 1598.673 -46.424

Antoine Equation:

log_{10}(P) = A − \frac{B}{T + C}

where P is the vapor pressure in bars and T is the temperature in K.

Compute the relative volatility at each point chosen to generate the x − y diagram. How close is each value as compared to the average value.